OpenStudy (anonymous):

What is the distance between (7, 1) and (-4, -5)?

OpenStudy (anonymous):

well there both in different quads 7,1 is in quad one and -4,-5 is in quad 3

OpenStudy (anonymous):

one is higher and one is lower

OpenStudy (anonymous):

one is further left and one is further right

OpenStudy (anonymous):

do thses replys help?

OpenStudy (anonymous):

well.. not really :/ here's the choices..\[3\sqrt{5} \]

OpenStudy (anonymous):

\[\sqrt{137}\]

OpenStudy (anonymous):

\[\sqrt{157}\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

those are all the choices and i still dont get it